Can you please show me how to get the answer for
3A divided by 4 + 1 = 2A-1 divided by 3.
I'm struggling with it..Thanks!

Answer by stanbon(75887)

3A divided by 4 + 1 = 2A-1 divided by 3.
------
[(3A)/4] + 1 = (2A-1)/3
-----
Multiply thru by 12 to get:
9A + 12 = 4(2A-1)
9A + 12 = 8A - 4
A = -16
-----
Cheers,
Stan H.